# Prediction Instability in Machine Learning Ensembles

### v1

@@ Abstract
In machine learning ensembles predictions from multiple models are aggregated. Despite widespread use and strong performance of ensembles in applied problems little is known about the mathematical properties of aggregating models and associated consequences for safe, explainable use of such models. In this paper we prove a theorem that shows that any ensemble will exhibit at least one of the following forms of prediction instability. It will either ignore agreement among all underlying models, change its mind when none of the underlying models have done so, or be manipulable through inclusion or exclusion of options it would never actually predict. As a consequence, ensemble aggregation procedures will always need to balance the benefits of information use against the risk of these prediction instabilities. This analysis also sheds light on what specific forms of prediction instability to expect from particular ensemble algorithms; for example popular tree ensembles like random forest, or xgboost will violate basic, intuitive monotonicity and fairness properties.
footnotetext: 1Department of Electrical Engineering and Computer Science, Milwaukee School of Engineering, Milwaukee, WI, USA. Correspondence to: Jeremy Kedziora <kedziora@msoe.edu>.
Preliminary work, Copyright 2024 by the author(s).
1 Introduction
@@
To motivate the axiomatic approach, consider a tree ensemble used for classification trained via e.g. the random forest algorithm [Breiman, 1996, 1999, 2001] or the xgboost algorithm [Chen and Guestrin, 2016]–a collection of $m$ decision trees with tree $j$ represented as a pair $(T_{j},L_{j})$ where $T_{j}$ is a partition of the data and $L_{j}(x)$ is a function mapping from an observation in the feature space $x$ to an element of the partition. One method to use this model to make predictions or choices would be to aggregate scores from the $m$ trees and choose the class $y$ with the largest such score:
@@
where $I(\cdot)$ is the indicator function. As an example, suppose that we fit such an ensemble on a data set with three classes and then, for a pair of points in the feature space $x_{1}$ and $x_{2}$, conduct inference to obtain the following:
@@
From this we would predict or choose class $1$ for $x_{1}$ and class 2 for $x_{2}$. But notice that these two points in feature space are similar in the sense that trees 2 and 3 have identical scores, and tree 1 yields that class 1 is more likely than class 2, which is more likely than class 3 in both cases. Despite this, after aggregation the ensemble prediction or choice is not the same for these two examples. That is, the ensemble “changed its mind” about the relative likelihood of the classes even though no individual model did.
@@
In this note we argue that this sort of prediction or decision-making inconsistency is an intrinsic part of ensembling. Specifically, we prove a theorem that shows that any ensemble will exhibit at least one of the following forms of prediction instability. It will either ignore agreement among all underlying models, change its mind when none of the underlying models have done so, or be manipulable through inclusion or exclusion of options it would never actually predict. As a consequence, ensemble aggregation procedures will always need to balance the benefits of information use against the risk of choice instability. We formalize this tradeoff in what we will call an ensemble choice discontinuity theorem.
@@
The remainder of this section lays out other work that we draw upon, the second section introduces our framework and axioms, the third specifies our main results, and the fourth discusses implications of these results and concludes.
Related Work
@@
The argument for ensembling many models together can be traced back to the jury theorems inspired by de Condorcet [1785]. These methods are heavily used in classical supervised learning [e.g. Breiman, 2001], but also appear in modern deep learning [Ganaie et al., 2022], in the token generation of large language models [e.g. Shazeer et al., 2017], and in reinforcement learning [Song et al., 2023].
@@
Modern use of ensembling techniques tends to involve choosing a base algorithm to construct learners, sampling data/features to train that base algorithm via bagging [Breiman, 1996, 1999, 2001] or boosting [Freund and Schapire, 1997] numerous times, and finally aggregating the outputs of individual models, e.g via hard or soft voting, or some learned procedure, e.g. stacking [Wolpert, 1992]. [^1] See Int [2000], Maclin and Opitz [2011], Polikar [2006], or Rokach [2010] for reviews. Despite the availability of numerous options for aggregation little scholarship has focused on studying the qualities of different aggregation procedures. As others [e.g. Wolpert and Macready, 1997, Werbin-Ofir et al., 2019] have noted, the optimal choice of aggregator is likely to be problem-specific. Given this, we study aggregators from an axiomatic perspective. This allows us to identify combinations of desirable properties, seek the aggregators that satisfy them, and thereby characterize the consistency and stability of the predictions or decision-making derived from ensembles.
@@
Social choice theory has long analyzed aggregative mechanisms axiomatically in the context of economics and political science. One of the earliest social choice results, Arrow’s impossibility theorem argued that no method of aggregating the ordinal preferences of a set of individuals into a single group-level ordinal preference could satisfy a set of desirable properties chosen to maximize consistency in decision-making and democratic principles [see Arrow, 1951].
@@
We take advantage of three specific extensions to Arrow’s original result to characterize ensemble aggregation. [^2] For a review see Patty and Penn [2019]. First, Arrow analyzed an orders-to-order mapping and machine learning models typically output scores, implying a scores-to-score mapping. Harsanyi [1955], Samuelson [1967], Sen [1970a], Schwartz [1970], Fishburn [1972], Kalai and Schmeidler [1977] and others extended Arrow’s result to cardinal preferences and we utilize their insights here. Second, in operational contexts machine learning is often used to make decisions, implying a focus on the choice derived from the scores-to-score mapping rather than the output score itself. This mirrors efforts to analyze the compatibility of Arrow’s axiomatic approach with so-called social choice functions that identify a subset of the alternatives rather than a preference ordering over them [e.g. Sen, 1969, 1970b, 1971] and that inform our own formalization. Finally, efforts have been made to show that Arrow’s axioms are special cases of other properties, leading to more general impossibility theorems. We will make use of Eliaz [2004] in this regard.
2 Framework
@@
With this response variable we will associate a set of predictor variables referred to as features and assumed to be an element $x$ of some feature data space $X$. Given a data set of examples training a model $j$ is equivalent to choosing a scoring function $s_{j}:X\to S_{j}(Y)\subseteq\mathbb{R}^{|Y|}$ where $s_{j}(x)=\{s_{j}(x,y)\}_{y\in Y}$. Beyond that, let $\mathbf{R}$ denote the set of all complete and transitive binary relations on $Y$; any of these binary relations will induce an ordering of the label values in $Y$. Given a particular ordering $R\in\mathbf{R}$ we can collect all the different scores consistent with this ordering as a set:
$F(Y,R)=\left\{z\in\mathbb{R}^{|Y|}\left|\mbox{for all }y,y^{\prime}\in Y\atop{y\hskip 2.84526ptR\hskip 2.84526pty^{\prime}\mbox{ iff }z_{d(y)}\geq z_{d(y^{\prime})}}\right.\right\}$
where $d(y)$ is the dimension associated with the score for $y$. We assume throughout that for all $R\in\mathbf{R}$ we have $S_{j}(Y)\bigcap F(Y,R)\neq\emptyset$ so that the model is not precluded from encoding any ordering of the label values. [^3] Note that this is not the same as assuming surjectivity of $s_{j}(\cdot)$. We will have more to say about how this relates to our results in the discussion section.
This model scoring function maps the feature inputs to scores that can be used to predict values of $y$ accurately, according to some well-defined notion of accuracy. In analogue with our motivating tree ensemble example, a model scoring function could be the output of a single decision tree.
Model Ensembles
@@
$s(x)=$$\left\{\left\{s_{j}(x,y)\right\}_{y\in Y}\right\}_{j=1,\ldots,m}$
$\hskip 56.9055pt\in\prod_{j}S_{j}(Y)\equiv S(Y).$
In general, we will suppress dependence on $x$ unless we need to distinguish between different input features.
With a set of model scoring functions in hand we will model the ensemble prediction as a choice from the set of label values, $Y$. Accordingly, define an ensemble choice aggregator as a set-valued function: $C:Y\times S(Y)\rightarrow\mathcal{P}(Y)\backslash\emptyset$, where $\mathcal{P}(Y)$ is the power set of $Y$ [^4] We exclude $\emptyset$ from the range of $C(\cdot,s)$ to focus throughout on ensemble choice aggregators that actually make choices. Later, we will use this to rule out a form of intransitivity in Lemma 3.2. The job of this function is to use the model scores to make a choice of one or more options from $Y$. In general, we will assume for any $Y^{*}\subseteq Y$ and any $s$ that $y\in C(Y^{*},s)$ if there does not exist $y^{\prime}\in Y^{*}$ such that $y^{\prime}=C(\{y,y^{\prime}\},s)$.
Aggregator Properties
@@
    How do the decisions made by an ensemble respond to post-training changes to the set of labels the response could take on?
@@
    How the decisions made by an ensemble respond when we impose requirements on how the score aggregator uses information from the underlying models?
@@
Change to the set of possible labels comes in two forms: deletions and insertions. For example, an online retailer using a recommender system could choose to permanently drop products from its catalogue; if it is expensive to retrain models then deleting a label option would be a temporary method to ensure that recommendations remain in sync with available products. We propose the following definition–adapted to our context from Arrow [1959] and Sen [1969]–to formalize ensemble decision consistency given these types of changes:
Definition 2.1.
@@
Say that $C(\cdot,s)$ is insertion/deletion consistent if for all $s\in S(Y)$ and any $Y_{1}^{*},Y_{2}^{*}\subseteq Y$ such that $Y_{1}^{*}\subseteq Y_{2}^{*}$ we have that $Y_{1}^{*}\bigcap C(Y_{2}^{*},s)\neq\emptyset$ implies that $C(Y_{1}^{*},s)=Y_{1}^{*}\bigcap C(Y_{2}^{*},s)$.
@@
Insertion/deletion consistency requires that if the ensemble chooses or predicts some label options when all possible labels for the response are available and then the set of labels is reduced but still contains some previously chosen labels then 1) no label passed over before now becomes chosen and 2) no previously chosen label is now passed over. In other words, the ensemble cannot be compelled to change what it does choose via requiring that it ignore label options it never chose in the first place; adding new label options cannot break ties. [^5] Sen [1969] gives the following examples of this property for intuition on page 384 of his original paper: “…if the world champion in some game is a Pakistani, then he must also be the champion in Pakistan…if some Pakistani is a world champion, then all champions of Pakistan must be champions of the world.”
@@
We turn next to consider desirable requirements on how an ensemble uses information, adapted from Arrow [1951], Sen [1970a], and Eliaz [2004].
Definition 2.2.
@@
Say that $C(\cdot,s)$ is a nondegenerate ensemble choice aggregator if there does not exist $j$ such that for all $y,y^{\prime}\in Y$ and all $s\in S(Y)$ if $s_{j}(\cdot,y)>s_{j}(\cdot,y^{\prime})$ then $y=C(\{y,y^{\prime}\},s)$.
Definition 2.3.
@@
Say that $C(\cdot,s)$ satisfies ensemble unanimity if for all $y,y^{\prime}\in Y$ and all $s\in S(Y)$ we have that $s_{j}(\cdot,y)>s_{j}(\cdot,y^{\prime})$ for all $j$ implies $y=C(\{y,y^{\prime}\},s)$.
Definition 2.4.
@@
Say that $C(\cdot,s)$ respects model choice reversal if for all $y,y^{\prime}\in Y^{*}\subseteq Y$ and all $s,s^{\prime}\in S(Y)$ we have that $y\in C(\{y,y^{\prime}\},s)$, $y^{\prime}\notin C(\{y,y^{\prime}\},s)$, and $y^{\prime}\in C(\{y,y^{\prime}\},s^{\prime})$ implies that there exists $j$ such that $s_{j}(\cdot,y)>s_{j}(\cdot,y^{\prime})$ and $s^{\prime}_{j}(\cdot,y)<s^{\prime}_{j}(\cdot,y^{\prime})$.
@@
First, if the intuition of ensembling–that many models improve on decision-making over a single model–is born out then the ensemble should use the information of multiple models to produce its decisions and the aggregator should not be an isomorphism of any single model. Definition 2.2 is a minimal version of such a requirement in that it requires that a single model does not completely determine the final choice. Second, it seems natural to assume that if every model in the ensemble scores $y$ higher than $y^{\prime}$ then the choice of the ensemble should reflect this; we formalize this in Definition 2.3. Third, how should the propensity of the ensemble to change its decision between labels depend on the propensity of the underlying models to do so? In Definition 2.4 we stipulate that the ensemble cannot change its mind for $y$ and $y^{\prime}$ unless at least one model in the ensemble changed its mind for $y$ and $y^{\prime}$. [^6] Model choice reversal is a generalization of the independence of irrelevant alternatives property used by Arrow [1951] and many others. See Eliaz [2004] Propositions 1 and 2 for analysis.
3 An Ensemble Choice Discontinuity Theorem

### v5

@@ Abstract
In machine learning ensembles predictions from multiple models are aggregated. Despite widespread use and strong performance of ensembles in applied problems little is known about the mathematical properties of aggregating models and associated consequences for safe, explainable use of such models. In this paper we prove a theorem that shows that any ensemble will exhibit at least one of the following forms of prediction instability. It will either ignore agreement among all underlying models, change its mind when none of the underlying models have done so, or be manipulable through inclusion or exclusion of options it would never actually predict. As a consequence, ensemble aggregation procedures will always need to balance the benefits of information use against the risk of these prediction instabilities. This analysis also sheds light on what specific forms of prediction instability to expect from particular ensemble algorithms; for example popular tree ensembles like random forest, or xgboost will violate basic, intuitive fairness properties. Finally, we show that this can be ameliorated by using consistent models in asymptotic conditions.
Index Terms:
machine learning ensembles, statistical learning theory, social choice theory
I Introduction
@@
To motivate the axiomatic approach, consider a tree ensemble used for classification trained via e.g. the random forest algorithm [4, 5, 6] or the xgboost algorithm [7]–a collection of $m$ decision trees with tree $j$ represented as a pair $(T_{j},L_{j})$ where $T_{j}$ is a partition of the data and $L_{j}(x)$ is a function mapping from an observation in the feature space $x$ to an element of the partition. One method to use this model to make predictions or choices would be to aggregate scores from the $m$ trees and choose the class $y$ with the largest such score:
@@
where $I(\cdot)$ is the indicator function. As an example, suppose that we fit such an ensemble on a data set with three classes and then, for a pair of points in the feature space $x_{1}$ and $x_{2}$, conduct inference to obtain the results in Table I.
Table I:
@@
From Table I we would predict or choose class $1$ for $x_{1}$ and class 2 for $x_{2}$. But notice that these two points in feature space are similar in the sense that trees 2 and 3 have identical scores, and tree 1 yields that class 1 is more likely than class 2, which is more likely than class 3 in both cases. Despite this, after aggregation the ensemble prediction or choice is not the same for these two examples. That is, the ensemble “changed its mind” about the relative likelihood of the classes even though no individual model did.
@@
In this note we argue that this sort of prediction or decision-making inconsistency is an intrinsic part of ensembling. Specifically, we prove a theorem that shows that any ensemble will exhibit at least one of the following forms of prediction instability. It will either ignore agreement among all underlying models, change its mind when none of the underlying models have done so, or be manipulable through inclusion or exclusion of options it would never actually predict. As a consequence, ensemble aggregation procedures will always need to balance the benefits of information use against the risk of choice instability. This result is driven by the capacity of the individual models within the ensemble and so provides an additional axiomatic rationale for the use of weak learners within an ensemble. We also show that this prediction instability is eliminated when using consistent models in asymptotic conditions and so can be thought of as a finite-sample property of ensembling.
@@
The remainder of this section lays out other work that we draw upon, the second section introduces our framework and axioms, the third section specifies our result on ensemble prediction instability, the fourth section discusses implications of these results, and the fifth section establishes our asymptotic result.
I-A Related Work
@@
The argument for ensembling many models together can be traced back to the jury theorems inspired by de Condorcet [8]. These methods are heavily used in classical supervised learning [e.g. 6], but also appear in modern deep learning [13], in the token generation of large language models [e.g. 28], and in reinforcement learning [29].
@@
Modern use of ensembling techniques tends to involve choosing a base algorithm to construct learners, sampling data/features to train that base algorithm via bagging [4, 5, 6] or boosting [12] numerous times, and finally aggregating the outputs of individual models, e.g via hard or soft voting, or some learned procedure, e.g. stacking [32]. [^1] See [16], [18], [20], or [21] for reviews. Despite the availability of numerous options for aggregation little scholarship has focused on studying the qualities of different aggregation procedures. As others [e.g. 31, 30] have noted, the optimal choice of aggregator is likely to be problem-specific. Given this, we study aggregators from an axiomatic perspective. This allows us to identify combinations of desirable properties, seek the aggregators that satisfy them, and thereby characterize the consistency and stability of the predictions or decision-making derived from ensembles.
@@
Social choice theory has long analyzed aggregative mechanisms axiomatically in the context of economics and political science. One of the earliest social choice results, Arrow’s impossibility theorem argued that no method of aggregating the ordinal preferences of a set of individuals into a single group-level ordinal preference could satisfy a set of desirable properties chosen to maximize consistency in decision-making and democratic principles [see 1].
@@
We take advantage of three specific extensions to Arrow’s original result to characterize ensemble aggregation. [^2] For a review see [19]. First, Arrow analyzed an orders-to-order mapping and machine learning models typically output scores, implying a scores-to-score mapping. [15], [22], [25], [23], [11], [17] and others extended Arrow’s result to cardinal preferences and we utilize their insights here. Second, in operational contexts machine learning is often used to make decisions, implying a focus on the choice derived from the scores-to-score mapping rather than the output score itself. This mirrors efforts to analyze the compatibility of Arrow’s axiomatic approach with so-called social choice functions that identify a subset of the alternatives rather than a preference ordering over them [e.g. 24, 26, 27] and that inform our own formalization. Finally, efforts have been made to show that Arrow’s axioms are special cases of other properties, leading to more general impossibility theorems. We will make use of [9] in this regard.
II Framework
@@
With this response variable we will associate a set of predictor variables referred to as features and assumed to be an element $x$ of some feature data space $X$. Given a data set of examples $D_{n}=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ training a model $j$ is equivalent to choosing a scoring function $s_{j}:X\to S_{j}(Y)\subseteq\mathbb{R}^{|Y|}$ where $s_{j}(x)=\{s_{j}(x,y)\}_{y\in Y}$. When we need to express the dependence of $s_{j}(\cdot)$ on the training data $D_{n}$ we will write $s_{j}(x|D_{n})$; otherwise, we will simply write $s_{j}(x)$. This model scoring function maps the feature inputs to scores that can be used to predict values of $y$ correctly, according to some well-defined measure of correctness. In analogue with our motivating tree ensemble example, a model scoring function could be the output of a single decision tree.
II-A Model Ensembles
@@
$s(x)=$$\left\{\left\{s_{j}(x,y)\right\}_{y\in Y}\right\}_{j=1,\ldots,m}\in\prod_{j}S_{j}(Y)\equiv S(Y).$
We will write $s(x)$ to mean the function mapping from $X$ to $S(Y)$ as defined above. When we want to work with an arbitrary element of $S(Y)$ rather than the function we will write it as $z=\{\{z_{j,d(y)}\}_{y\in Y}\}_{j=1,\ldots,m}$ where $d(y)$ is the dimension measuring the score for $y$, and where $s(x)=z$ if and only if $s_{j}(x,y)=z_{j,d(y)}$ for all $j$ and all $y\in Y$.
With a set of model scoring functions in hand we will model the ensemble prediction as a choice from the set of label values, $Y$. Accordingly, we formalize the method for using the scores from many models to make predictions or choices in the following way. [^3] We exclude $\emptyset$ from the range of $C(\cdot)$ to focus throughout on ensemble choice aggregators that actually make choices. Later, we will use this to rule out a form of intransitivity in Lemma III.2.
Definition II.1.
An ensemble choice aggregator is a set-valued function $C:2^{Y}\backslash\{\emptyset\}\times S(Y)\rightarrow 2^{Y}\backslash\{\emptyset\}$ such that for any $Y^{*}\subseteq Y$ and any $z\in S(Y)$ we have $C(Y^{*},z)\subseteq Y^{*}$ and also that $y\in C(Y^{*},z)$ if there does not exist $y^{\prime}\in Y^{*}$ such that $\{y^{\prime}\}=C(\{y,y^{\prime}\},z)$.
The job of this function is to use the model scores to make a choice of one or more options from a subset of $Y$ [^4] The ensemble choice aggregator is purposefully abstract and general to maximize the reach of our theoretical analysis. A common, specific ensemble choice aggregator would be to choose the response label that maximizes the average score across models given the input. We will analyze this below. Finally, an ensemble is a pair $(C,s)$ consisting of the set of individual model scoring functions and the method of aggregating them.
II-B Model Capacity Properties
The representational capacity of the individual models in an ensemble will turn out to be important in our results. We will use two properties to capture model capacity. First, the extent to which scores consistent with each possible ordering of the response labels are available in the range of an individual model bounds the flexibility of that model; each potential ordering of response labels excluded from $S_{j}(Y)$ would prevent the model from learning that ordering no matter what information is available in the data.
To formalize this idea, let $\mathbf{R}$ denote the set of all complete and transitive binary relations on $Y$; any of these binary relations will induce an ordering of the response labels in $Y$. Given a particular weak ordering $R\in\mathbf{R}$ we can collect all the different scores consistent with this ordering as a set:
$W(Y,R)=\left\{z\in\mathbb{R}^{|Y|}\left|\mbox{for all }y,y^{\prime}\in Y\atop{y\hskip 2.84526ptR\hskip 2.84526pty^{\prime}\mbox{ iff }z_{d(y)}\geq z_{d(y^{\prime})}}\right.\right\}$
A potentially high capacity model would then be one where no ordering of label values is excluded from the model’s range. Formally:
Definition II.2.
Say that model $j$ is an unconstrained learner if for all $R\in\mathbf{R}$ we have $S_{j}(Y)\bigcap W(Y,R)\neq\emptyset$.
An unconstrained learner is one in which scores could be learned that correspond to any ordering of the response labels.
Second, we can also be rigorous about exactly how many of these orderings a model actually learns. Formally:
Definition II.3.
Given $s(x)$ as defined above say that $S^{*}\subseteq S(Y)$ is observable by $\mathbf{s(x)}$ if for all $z\in S^{*}$ there exists $x\in X$ such that $s(x)=z$.
Under this definition, a given ordering over the possible label values is learned by a model if there exists a point in the feature space that maps to scores that correspond to it. [^5] Note neither of these definitions are the same as assuming surjectivity of $s_{j}(\cdot)$. We will have more to say about how this relates to our results below.
II-C Aggregator Properties
@@
    How do the predictions or choices made by an ensemble respond to post-training changes to the set of labels the response could take on?
@@
    How the predictions or choices made by an ensemble respond when we impose requirements on how the score aggregator uses information from the underlying models?
@@
Change to the set of possible labels comes in two forms: deletions and insertions. For example, an online retailer using a recommender system could choose to permanently drop products from its catalogue; if it is expensive to retrain models then deleting a label option would be a temporary method to ensure that recommendations remain in sync with available products. We propose the following definition–adapted to our context from [2] and [24]–to formalize ensemble decision consistency given these types of changes:
Definition II.4.
@@
Say that $C(\cdot)$ is insertion/deletion consistent if, for all $z\in S(Y)$ and all $Y_{1}^{*},Y_{2}^{*}\subseteq Y$ such that $Y_{1}^{*}\subseteq Y_{2}^{*}$, we have that $Y_{1}^{*}\bigcap C(Y_{2}^{*},z)\neq\emptyset$ implies that $C(Y_{1}^{*},z)=Y_{1}^{*}\bigcap C(Y_{2}^{*},z)$.
@@
Insertion/deletion consistency requires that if the ensemble chooses or predicts some label options when all possible labels for the response are available and then the set of labels is reduced but still contains some previously chosen labels then 1) no label not chosen before now becomes chosen and 2) no previously chosen label is now not chosen. In other words, the ensemble cannot be compelled to change what it does choose via requiring that it ignore label options it never chose in the first place; adding new label options cannot break ties.666[24] gives the following examples of this property for intuition on page 384 of his original paper: “…if the world champion in some game is a Pakistani, then he must also be the champion in Pakistan…if some Pakistani is a world champion, then all champions of Pakistan must be champions of the world.”
@@
We turn next to consider desirable requirements on how an ensemble uses information, adapted from Arrow [1], Sen [25], and Eliaz [9].
Definition II.5.
@@
Say that $C(\cdot)$ is a nondegenerate ensemble choice aggregator if there does not exist $j$ such that, for all $y,y^{\prime}\in Y$ and all $z\in S(Y)$, we have that $z_{j,d(y)}>z_{j,d(y^{\prime})}$ implies that $\{y\}=C(\{y,y^{\prime}\},z)$.
Definition II.6.
@@
Say that $C(\cdot)$ satisfies ensemble unanimity if, for all $y,y^{\prime}\in Y$ and all $z\in S(Y)$, we have that $z_{j,d(y)}>z_{j,d(y^{\prime})}$ for all $j$ implies $\{y\}=C(\{y,y^{\prime}\},z)$.
Definition II.7.
@@
Say that $C(\cdot)$ respects model choice reversal if, for all $y,y^{\prime}\in Y^{*}\subseteq Y$ and all $z,z^{\prime}\in S(Y)$, we have that $y\in C(\{y,y^{\prime}\},z)$, $y^{\prime}\notin C(\{y,y^{\prime}\},z)$, and $y^{\prime}\in C(\{y,y^{\prime}\},z^{\prime})$ implies that there exists $j$ such that $z_{j,d(y)}>z_{j,d(y^{\prime})}$ and $z^{\prime}_{j,d(y)}<z^{\prime}_{j,d(y^{\prime})}$.
@@
First, if the intuition of ensembling–that many models improve on decision-making over a single model–is born out then the ensemble should use the information of multiple models to produce its decisions and the aggregator should not be an isomorphism of any single model. Definition II.5 is a minimal version of such a requirement in that it requires that a single model does not completely determine the final choice. Second, it seems natural to assume that if every model in the ensemble scores $y$ higher than $y^{\prime}$ then the choice of the ensemble should reflect this; we formalize this in Definition II.6. Third, how should the propensity of the ensemble to change its decision between labels depend on the propensity of the underlying models to do so? In Definition II.7 we stipulate that the ensemble cannot change its mind for $y$ and $y^{\prime}$ unless at least one model in the ensemble changed its mind for $y$ and $y^{\prime}$. [^7] Model choice reversal is a generalization of the independence of irrelevant alternatives property used by [1] and many others. See [9] Propositions 1 and 2 for analysis.
III An Ensemble Aggregated–Prediction Instability Theorem